OpenStudy (ksaimouli):

how to graph this with in 30 sec

OpenStudy (ksaimouli):



OpenStudy (ksaimouli):

this is deferentiable or not how can i say /

OpenStudy (ksaimouli):

@Callisto

OpenStudy (ksaimouli):

how can i figure out whether this function is differentiable or not

OpenStudy (zepp):

This function is differentiable only at where this function exists

OpenStudy (zepp):

The derivative of this function would be \[\large \frac{d}{dx}(x^{\frac{2}{5}})=\frac{2}{5x^{\frac{3}{5}}}\]

OpenStudy (zepp):

Differentiable from 0 to \(+\infty\)

OpenStudy (zepp):

Wait, I'm not sure :/

OpenStudy (callisto):

Hmm... A function is differentiable when the left hand limit = right hand limit as x-> c, I think? Of course, that function must be continuous. Perhaps you can read this: http://www.math.wfu.edu/tutorials/Math111/ContinuityandDifferentiability.pdf Check the part differentiability.

OpenStudy (ksaimouli):

nope can any one explain how to graph without taking points

OpenStudy (ksaimouli):

this one has corner so not differentiated